In a class of 16 students, 13 students took a Math test, 10 students took Science test, and 9 students took both tests.
How many students did not take any of the tests?

Found 2 solutions by MathLover1, ikleyn:
Answer by MathLover1(20850)


In a class of 16 students,
if 13 students took a Math test, so far 3 students did not take a Math test
if 10 students took Science test, and 9 students took both tests, then there was 1 student who took Science test only
so, if 13 students took a Math test and 1 student who took Science test only, means 14students took either a Math test or Science test or both
so 16-14=2 students did not take any of the tests

Answer by ikleyn(52900)

.
In a class of 16 students, 13 students took a Math test, 10 students took Science test, and 9 students took both tests.
How many students did not take any of the tests?
--------------------------------------------------------

How many students took the Math test only and didn't took the Science test?
Their number is 13 - 9 = 4.

How many students took the Science test only and didn't took the Math test?
Their number is 10 - 9 = 1.

How many students took at least one, the Math or the Science test?
Their number is 9 + 4 + 1 = 14.

Hence, 16-14 = 2 students didn't take any of the tests.

Answer. 2.